A price-taking firm in a perfectly competitive market faces a market price of $4. The firm's marginal cost function is MC(Q) = 2 + aQ, where "a" is a positive number. As "a" increases, the firm's profit-maximizing quantity increases, decreases, or does not change?
Answer #1

Answer

decrease

=====

the frim produces at MC=P

2+aQ=4

aQ=2

Q=2/a

the 'a' is inversely related to the quantity so an increase in 'a' decreases the quantity.

Ex

a=1

Q=2/1=2

a=2

Q=2/2=1

so the Q decreases as a increases.